an = a 0.5n + n, a1 = 0, where n is a power of 2, is a linear recurrence relation. ( true / falsa ).


1
Expert's answer
2021-05-17T13:50:02-0400

A linear recurrence relation is an equation that expresses each element of a sequence as a linear function of the preceding ones. More precisely, in the case where only the immediately preceding element is involved, a linear recurrence relation has the form "u_n=\varphi (n,u_{n-1},u_{n-2},\ldots ,u_{n-d})," where "{\displaystyle \varphi :\mathbb {N} \times X^{d}\to X} , \ \varphi (n,u_{n-1},u_{n-2},\ldots ,u_{n-d})=c_1u_{n-1}+c_2u_{n-2}+\ldots +c_da_{n-d}+f(n),"

is a everywhere defined function that involves "d"  consecutive elements of the sequence. In this case, "d"  initial values are needed for defining a sequence.


Taking into account that "a_n = a_{0.5n} + n" is defined only for "n" equals to the powers of 2, and hence is not defined for the rest natural numbers, the previous definition implies that "a_n = a_{0.5n} + n" is not a linear recurrence relation.


Answer: false